OpenStudy (konradzuse):

Solve the differential Equation dy/dx = (8e^x - 9e^-x)^2

OpenStudy (aliahmed):

please 8^x or x^8 ????

OpenStudy (konradzuse):

err crap (8e^x - 9e^-x)^2

OpenStudy (aliahmed):

2(8e^x-9e^-x)^1*8xe^x+9xe^-x

terenzreignz (terenzreignz):

@aliahmed You differentiated, but I think we needed to integrate :)

OpenStudy (aliahmed):

why?

OpenStudy (turingtest):

...so expand, separate the variables, and integrate

OpenStudy (turingtest):

because we want y and are given dy/dx

terenzreignz (terenzreignz):

The entire equation was equal to dy/dx, so we have to reverse the process, don't we?

terenzreignz (terenzreignz):

sorry, not equation, but expression, my bad :)

OpenStudy (aliahmed):

yeye we want y

OpenStudy (konradzuse):

Apparently for (8e^x - 4e^-x)^2 is 32e^(2x) - 64x -8^(-2x) +c

OpenStudy (turingtest):

is it a 9 or a 4 for the other coefficient?

OpenStudy (konradzuse):

it's a 9, but this was a practice question I was doing with a 4.

OpenStudy (konradzuse):

So when we expand 8^x * 9^-x does it become 63^-2x?

OpenStudy (konradzuse):

with the e's in there too :P

terenzreignz (terenzreignz):

Why did you multiply them? I thought they were to be subtracted?

OpenStudy (turingtest):

there is no x that comes out

OpenStudy (konradzuse):

(8e^x - 9e^-x)^2 wouldnt' we do (8e^x - 9e^-x) (8e^x - 9e^-x)????

OpenStudy (konradzuse):

and then cross multiply?

terenzreignz (terenzreignz):

well, expand (8e^x - 9e^-x)^2, whatever you call expanding it

OpenStudy (turingtest):

it would be what you perhaps know as FOIL ?

terenzreignz (terenzreignz):

Yes :)

OpenStudy (konradzuse):

yes

OpenStudy (turingtest):

so FOIL out the expression above

OpenStudy (turingtest):

...that is the same thing as "expanding" (8e^x-9e^-x)^2

OpenStudy (konradzuse):

64e^2x - 72e^-2x + 81e^2x?

OpenStudy (turingtest):

not quite; careful with those exponents...

terenzreignz (terenzreignz):

Not to mention you have to take the product of the two terms twice

OpenStudy (konradzuse):

yeah I was konfused with the exponents.

OpenStudy (turingtest):

what are the terms? look at the outer term it will be -(8e^x)(9e^-x) can you simplify that?

terenzreignz (terenzreignz):

By terms I meant 8e^x and 9e^-x ^^

OpenStudy (konradzuse):

oh woopsies I forgot to do the middle terms twice :P. Would it be -72e^-2x?

OpenStudy (konradzuse):

so 2x that would be 144e^4x?

terenzreignz (terenzreignz):

no, check your exponents again 8e^x 9e^-x :)

OpenStudy (turingtest):

your law of exponents is\[x^a\cdot x^b=x^{a+b}\]

OpenStudy (konradzuse):

Idk I suck at exponents! They are supposed to be multiplied right ?? X * -X?

OpenStudy (konradzuse):

oh okay I was thinking that too.. SO it wouldbe x-x? so 0?

OpenStudy (turingtest):

yes, because they have the same base we add exponents when we multply

OpenStudy (turingtest):

and the x+(-x)=0 so what is \[\large 9e^x\cdot8e^{-x}\]?

OpenStudy (konradzuse):

well e^0 =i 1 so would it just be 72?

OpenStudy (turingtest):

yes and the other term?

OpenStudy (konradzuse):

ugh where did that 'i' come from haha.

OpenStudy (turingtest):

I didn't even see it, lol

OpenStudy (konradzuse):

so it wouldbe -72-72 so -144?

OpenStudy (konradzuse):

:P

OpenStudy (turingtest):

yep :)

OpenStudy (konradzuse):

so then we get 64e^2x - 144 + 81^-2x?

OpenStudy (turingtest):

correct, and now we integrate

OpenStudy (turingtest):

rather pause that let's look at the process if it's new to you

OpenStudy (turingtest):

so far we have expanded the RHS to give\[\frac{dy}{dx}=64e^{2x}-144+81^{-2x}\]now we "separate the variables that is, we treat the symbol on the left as a fraction (though really, it isn't) and get the dx on the side with the x's like we would in algebra)

OpenStudy (turingtest):

\[\frac{dy}{dx}=64e^{2x}-144+81^{-2x}\]\[dy=64e^{2x}-144+81^{-2x}dx\]now we can integrate both sides\[\int dy=\int64e^{2x}-144+81^{-2x}dx\]what do you get?

OpenStudy (konradzuse):

well e^u is still e^u so those exponents stay the same.

OpenStudy (konradzuse):

but according to the practice version is seems as though you divide by 2 so would it be 32e^2x - 144 - 81^-2x/2?

OpenStudy (konradzuse):

I guess that's because our u = 2x. But actually in that case it should be +(81^-2x)/2

OpenStudy (konradzuse):

hmm guess that's not right./... It also says in the practice version answer that the middle term is *x? so 144x? Idk if that makes sense... :(

terenzreignz (terenzreignz):

Try doing it yourself, try to forget that you ever saw that practice version :)

OpenStudy (konradzuse):

I did as you can see above, but it's not right....

terenzreignz (terenzreignz):

Well, if you think it's not right, let's just start it again : \[\int\limits_{}^{}(64e^{2x} + 81e^{-x} - 144) dx\] How do you go about solving this integral?

OpenStudy (turingtest):

\[\int\limits_{}^{}(64e^{2x} + 81e^{-2x} - 144) dx\] ;)

OpenStudy (konradzuse):

^ yes lol

OpenStudy (konradzuse):

Well we can split it up into 3 integrals. u = 2x for the first, and u = -2x for the next.

OpenStudy (konradzuse):

it should come out to 32e^2x -81e^2x/2 - 144?

OpenStudy (konradzuse):

which is what I posted before...

OpenStudy (konradzuse):

or if u is just 2x then it would be +81e^2x/2

OpenStudy (turingtest):

you forgot integrate -144

OpenStudy (konradzuse):

oh yeah so 144x

OpenStudy (konradzuse):

Ok we win :). yay!

OpenStudy (konradzuse):

I always forget stupid stuff like that... Thanks for your help as always Turing!

OpenStudy (turingtest):

always welcome :)